# LANCE: Low Rank Activation Compression for Efficient On-Device Continual Learning

### v1

@@ 3.1 Results: Single-task Fine-Tuning Efficiency
We first evaluate LANCE on fine-tuning the last 2 or 4 layers of MCUNet (Lin et al., 2020), MobileNetV2 (Sandler et al., 2018), ResNet18 (He et al., 2016) and ResNet34 (He et al., 2016). All models are pretrained on ImageNet-1k (Krizhevsky et al., 2012). For LANCE, the offline Phase I calibration uses N=100N{=}100 mini-batches with an energy threshold $\varepsilon{=}0.7$. Fine-tuning runs for 50 epochs.
@@
Answer to Q2. Across all datasets and models, LANCE reduces activation storage by up to $250\times$ relative to BP while maintaining competitive accuracy. Table 1 reports results on CIFAR-10/100 and Pets and CUB200. Compared to iterative low-rank methods such as ASI (Nguyen et al., 2025), or HOSVD (Nguyen et al., 2024) where a full SVD is recomputed at each step, LANCE achieves comparable compression with higher accuracy and similar FLOPs, confirming that repeated decompositions are unnecessary. Although BP attains higher accuracy when the same number of layers are fine-tuned, LANCE’s extreme memory efficiency enables fine-tuning more layers under strict SRAM budgets. For example, fine-tuning 4 layers with LANCE reaches the accuracy of 2-layer BP while still fitting within microcontroller constraints. These results demonstrate that fixed low-rank subspaces retain sufficient representational power for efficient on-device fine-tuning.
@@
Table 2: Performance comparison on continual image classification datasets using multi-head networks. Accuracy and BWT (mean $\pm$ std) are reported over five trials. denotes the results taken from Saha et al. (2021) and denote the results from the respective original papers.
@@ 3.2 Results: Continual Learning
We compare LANCE against a broad set of continual learning baselines: regularization-based methods (EWC (Kirkpatrick et al., 2017), HAT (Serrà et al., 2018)), replay-based approaches (A-GEM (Chaudhry et al., 2019b), ER_Res (Chaudhry et al., 2019a)), and projection-based methods (GPM (Saha et al., 2021), TRGP (Lin et al., 2022b), CUBER (Lin et al., 2022c), SGP (Saha & Roy, 2023), CODE-CL (Apolinario et al., 2024)). For reference, we also report multitask performance, which serves as an upper bound where all tasks are trained jointly. We follow the same experimental setup as Apolinario et al. (2024): a 5-layer AlexNet for Split CIFAR-100, and a reduced ResNet18 for Split MiniImageNet and the 5-Datasets benchmark. During the first task, models are trained without constraints using full BP. For subsequent tasks, the model from the previous task serves as initialization, and LANCE applies its subspace-constrained training.
Answer to Q3. LANCE achieves accuracy competitive with full-rank gradient projection methods such as GPM and CODE-CL, while using drastically less memory. As summarized in Table 2, on Split CIFAR-100 LANCE reaches $\sim$71% average accuracy (ACC) while reducing activation storage by more than $10\times$ compared to GPM. Similar trends are observed on Split MiniImageNet and the 5-Datasets benchmark, where LANCE maintains stable performance without rehearsal memory. These results confirm that fixed low-rank subspaces are sufficient to retain representational capacity across tasks, while providing substantial memory savings for resource-constrained devices.
@@ 3.3 Ablations and Analysis
We conduct ablations on two key hyperparameters: (i) the energy threshold $\varepsilon$ (evaluated on CUB-200), and (ii) the number of calibration batches $N$ (evaluated on CIFAR-100 with $\varepsilon{=}0.7$). Results for MCUNet and ResNet34 are shown in Fig. 3. As expected, higher $\varepsilon$ increases accuracy but also grows memory usage exponentially, highlighting the need to balance compression and accuracy according to device memory budgets. For calibration size $N$, accuracy remains stable across a wide range, while memory usage grows with $N$. Remarkably, even N=2N{=}2 calibration batches are sufficient to yield stable subspaces, underscoring the robustness of the one-shot calibration phase.
Related Work.
@@
The memory cost of storing intermediate activations has long been recognized as the central bottleneck for training on resource-constrained devices (Cai et al., 2020; Lin et al., 2022a). Prior work has reduced this cost through recomputation (e.g., checkpointing (Chen et al., 2016)), architectural modifications (e.g., reversible networks (Gomez et al., 2017)), or lossy compression and sparsity (Chen et al., 2021; Barley & Fröning, 2023; Jiang et al., 2022), but often at the expense of extra computation, accuracy, or hardware complexity. A complementary line of work exploits low-rank structure in activations (Nguyen et al., 2024; 2025). Closest to our method is ASI (Nguyen et al., 2025), which incrementally updates subspaces during training; in contrast, LANCE performs a one-shot HOSVD at initialization and reuses the resulting subspaces throughout fine-tuning, avoiding repeated decompositions. This simple design not only reduces compute and hardware overhead but also enables a continual learning extension: fixed subspaces naturally support null-space allocation across tasks without storing large task-specific matrices. System-level methods such as TinyTL (Cai et al., 2020), MCUNet (Lin et al., 2020), and the 256KB training engine (Lin et al., 2022a) co-design architectures and update rules to fit extreme SRAM budgets, often by freezing layers or pruning gradient paths. LANCE is complementary to these approaches: instead of restricting the optimization, it preserves full backpropagation while directly compressing activations. Finally, in continual learning, regularization methods (Kirkpatrick et al., 2017; Serrà et al., 2018), replay-based approaches (Chaudhry et al., 2019b; a), and gradient-projection methods (Saha et al., 2021; Lin et al., 2022b; c; Saha & Roy, 2023; Apolinario et al., 2024) all mitigate forgetting by controlling parameter updates or storing exemplars. LANCE differs by operating directly in the activation subspace: new tasks are assigned to the orthogonal complement of previously used directions during calibration. This yields performance competitive with gradient-projection methods while requiring significantly less memory, and without rehearsal buffers.
@@ Acknowledgments
Appendix A Datasets, Models and Hyperparameters
@@
For single-task fine-tuning we report top-1 accuracy on the test split, peak activation memory in MB (estimated from tensors retained for the backward pass), and training FLOPs. [^1] Peak activation memory and FLOPs formulas are detailed in Appendix C. For continual learning experiments (Section 2.3), we additionally report ACC and BWT following standard CL definitions.
A.1 For Single-Task Fine-Tuning Experiments
A.1.1 Datasets
We evaluate on five image classification datasets spanning scale and difficulty: CIFAR-10 (Krizhevsky, 2009), CIFAR-100 (Krizhevsky, 2009), Oxford-IIIT Pets (Parkhi et al., 2012), Flowers-102 (Nilsback & Zisserman, 2008), and CUB-200 (Wah et al., 2023). We use the official training/test splits provided by each dataset and do not carve out an additional validation set. For all datasets, images are resized to match the native input resolution of the chosen architecture (e.g., $224\times 224$ for ImageNet-pretrained models), and standard data augmentation is applied consisting of random horizontal flipping.
A.1.2 Model Architectures
For single-task experiments we use four pretrained models: ResNet18, ResNet34 (He et al., 2016), MobileNetV2 (Sandler et al., 2018), and MCUNet (Lin et al., 2020). All of them are initialized from ImageNet-1k pretrained weights. For each model, we unfreeze the last $N\in\{2,4\}$ convolutional layers and fine-tune them while the rest of the network remains frozen.
A.1.3 Training Protocol
@@ A.1.4 Hyperparameters
A.2 For Continual Learning Experiments
@@
A.2.1 Model Architecture
We employ two neural network architectures in our experiments: an AlexNet-inspired model following Serrà et al. (2018), and a Reduced ResNet18 as described in Lopez-Paz & Ranzato (2017). The AlexNet-like network integrates batch normalization (BN) layers after every convolutional and fully connected layer except the final classifier. BN parameters are updated only during the first task and kept fixed thereafter. Its architecture consists of three convolutional layers with $64$, $128$, and $256$ filters, using kernel sizes of $4\times 4$, $3\times 3$, and $2\times 2$, respectively, each followed by a $2\times 2$ max-pooling operation. Two fully connected layers with $2048$ units each are appended. ReLU activations are applied throughout, and dropout is used with rates of $0.2$ for the first two layers and $0.5$ for the rest. The Reduced ResNet18 follows the setup in Saha et al. (2021). For Split miniImageNet tasks, its first convolutional layer uses stride $2$, whereas for the 5-Datasets benchmark, stride $1$ is used. Across all experiments, models are trained with cross-entropy loss.
A.2.2 Dataset Statistics
Tables 5 and 4 summarize the datasets employed in our continual learning benchmarks. We adopt the same train/test splits as in Saha et al. (2021); Lin et al. (2022b); Saha & Roy (2023); Apolinario et al. (2024). For the 5-Datasets benchmark, grayscale samples are duplicated across three channels to match RGB input requirements. All images are resized to $32\times 32$, yielding an effective input dimension of $3\times 32\times 32$.
A.2.3 Hyperparameters
@@ A.3 Compute resources
Appendix B Proofs for Section 2.4
@@
Since $\mathsf{P}^{(l)}$ is an orthogonal projector (A1),
@@
Thus $\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}$ is a descent direction unless it is zero. ∎
@@
for $\eta\leq 1/L$. Summation gives $\sum_{k}\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}^{2}<\infty$ and thus $\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\to 0$. Any limit point $\bm{\theta}$$_{\star}$ satisfies $\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{\star})=\mathbf{0}$. ∎
See 2.2
Proof.
From Theorem 2.2, $\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\to 0$. Then
@@
Take a limit point $\bm{\theta}$$_{\star}$. ∎
Remark B.1 (Why the leakage bound is reasonable).
@@
Appendix C Complexity Analysis of LANCE

### v2

@@ 3.1 Results: Single-task Fine-Tuning Efficiency
We first evaluate LANCE on fine-tuning the last 2 or 4 layers of MCUNet [Lin2020MCUNet:Devices], MobileNetV2 [mobilenetv2], ResNet18 [he2016deep] and ResNet34 [he2016deep]. All models are pretrained on ImageNet-1k. For LANCE, the offline Phase I calibration uses N=100N{=}100 mini-batches with an energy threshold $\varepsilon{=}0.7$. Fine-tuning runs for 50 epochs.
Table 1: ImageNet-scale fine-tuning on Split ImageNet-1k. We pretrain each model on 500 ImageNet classes (Split A) and fine-tune on the remaining 500 classes (Split B).
ResNet18 ResNet34
Method Acc $\uparrow$ MB $\downarrow$ TFLOPs $\downarrow$ Acc $\uparrow$ MB $\downarrow$ TFLOPs $\downarrow$
BP (2 layers) 60.70 24.50 0.059 60.77 24.50 0.059
BP (4 layers) 64.28 61.25 0.090 64.21 49.00 0.118
LANCE (2 layers) 59.18 7.12 0.048 58.59 4.49 0.046
LANCE (4 layers) 62.20 16.59 0.075 62.66 10.79 0.093
@@
Answer to Q2. Across all datasets and models, LANCE reduces activation storage by up to $250\times$ relative to BP while maintaining competitive accuracy. Table˜2 reports results on CIFAR-10/100 and Pets and CUB200. Compared to iterative low-rank methods such as ASI [nguyen2025beyond], or HOSVD [nguyen2024activation] where a full SVD is recomputed at each step, LANCE achieves comparable compression with higher accuracy and similar FLOPs, confirming that repeated decompositions are unnecessary. Although BP attains higher accuracy when the same number of layers are fine-tuned, LANCE’s extreme memory efficiency enables fine-tuning more layers under strict SRAM budgets. For example, fine-tuning 4 layers with LANCE reaches the accuracy of 2-layer BP while still fitting within microcontroller constraints. These results demonstrate that fixed low-rank subspaces retain sufficient representational power for efficient on-device fine-tuning.
@@
Scalability to ImageNet-Scale Fine-Tuning.
[…]
Table 3: Performance comparison on continual image classification datasets using multi-head networks. Accuracy and BWT (mean $\pm$ std) are reported over five trials. denotes the results taken from [Saha2021GradientLearning] and denote the results from the respective original papers.
@@ 3.2 Results: Continual Learning
We compare LANCE against a broad set of continual learning baselines: regularization based methods (EWC [Kirkpatrick2017OvercomingNetworks], HAT [Serra2018OvercomingTask]), replay-based approaches (A-GEM [Chaudhry2019EfficientA-GEM], ER_Res [Chaudhry2019OnLearning]), and projection-based methods (GPM [Saha2021GradientLearning], TRGP [Lin2022TRGP:Learning], CUBER [Lin22CUBER], SGP [Saha2023ContinualProjection], SD [SD2023], LMSP [lmsp2025], CODE-CL [apolinario2024code]). For reference, we also report multitask performance, which serves as an upper bound where all tasks are trained jointly. We follow the same experimental setup as [apolinario2024code]: a 5-layer AlexNet for Split CIFAR-100, and a reduced ResNet18 for Split MiniImageNet and the 5-Datasets benchmark. During the first task, models are trained without constraints using full BP. For subsequent tasks, the model from the previous task serves as initialization, and LANCE applies its subspace-constrained training.
Answer to Q3. LANCE achieves accuracy competitive with full-rank gradient projection methods such as GPM and CODE-CL, while using drastically less memory. As summarized in Tab.˜3, on Split CIFAR-100 LANCE reaches $\sim$71% average accuracy (ACC) while reducing activation storage by more than $10\times$ compared to GPM. Similar trends are observed on Split MiniImageNet and the 5-Datasets benchmark, where LANCE maintains stable performance without rehearsal memory. These results confirm that fixed low-rank subspaces are sufficient to retain representational capacity across tasks, while providing substantial memory savings for resource-constrained devices.
@@ 3.3 Ablations and Analysis
We conduct ablations on two key hyperparameters: (i) the energy threshold $\varepsilon$ (evaluated on CUB-200), and (ii) the number of calibration batches $N$ (evaluated on CIFAR-100 with $\varepsilon{=}0.7$). Results for MCUNet and ResNet34 are shown in Figure˜3. As expected, higher $\varepsilon$ increases accuracy but also grows memory usage exponentially, highlighting the need to balance compression and accuracy according to device memory budgets. For calibration size $N$, accuracy remains stable across a wide range, while memory usage grows with $N$. Remarkably, even N=2N{=}2 calibration batches are sufficient to yield stable subspaces, underscoring the robustness of the one-shot calibration phase.
Figure 4: Gradient fidelity across compression ratios. We vary the energy threshold $\varepsilon$ controlling the effective memory compression of LANCE and report the angle (mean $\pm$ std) between full-BP gradients and LANCE-projected gradients over the final 10 epochs (out of 50). As $\varepsilon$ decreases, compression becomes more aggressive and gradient alignment degrades accordingly. Nevertheless, even under memory reductions of up to two orders of magnitude, LANCE maintains gradient directions within $\sim 70^{\circ}$ of full BP, indicating preserved descent directions despite strong compression.
Figure 5: End-to-end fine-tuning latency on a Raspberry Pi 3B+ (MCUNet on CIFAR-10, batch size 128). Averaged over five trials, LANCE consistently achieves the lowest forward, backward, and total training times across all layer depths.
We additionally evaluate how different compression ratios affect gradient fidelity by measuring the angle between full-BP gradients and LANCE-projected gradients across several values of $\varepsilon$, shown in Figure˜4. As expected, more aggressive compression (lower $\varepsilon$) leads to larger angles, but the degradation is smooth and controlled. Even at the strongest compression settings, reducing activation storage by up to two orders of magnitude, LANCE maintains gradient alignment within $\sim 70^{\circ}$ of full BP, indicating that the one-shot subspace preserves the dominant descent directions required for stable optimization.
3.4 Runtime Validation on Edge Hardware
[…]
4 Related Work
@@
Storing intermediate activations is the primary memory bottleneck for on-device training [lin2020tinytl, lin2022ondevice]. Prior work has reduced this cost through recomputation (e.g., checkpointing [chen2016checkpoint]), architectural modifications (e.g., reversible networks [gomez2017reversible]), or lossy compression and sparsity [chen2021actnn, barley2023activation, yu2022backrazor], but often at the expense of extra computation, accuracy, or hardware complexity. A complementary line of work exploits low-rank structure in activations [nguyen2024activation, nguyen2025beyond]. Closest to our method is ASI [nguyen2025beyond], which incrementally updates subspaces during training; in contrast, LANCE performs a one-shot HOSVD at initialization and reuses the resulting subspaces throughout fine-tuning, avoiding repeated decompositions. Recent CL work such as LMSP [lmsp2025] also uses low-rank representations but focuses on reducing the computational cost of gradient projection, still requiring multiple per-task low-rank memory components; by contrast, LANCE targets the dominant activation-memory bottleneck directly. This simple design not only reduces compute and hardware overhead but also enables a continual learning extension: fixed subspaces naturally support null-space allocation across tasks without storing large task-specific matrices. System-level methods such as TinyTL [lin2020tinytl], MCUNet [Lin2020MCUNet:Devices], and the 256KB training engine [lin2022ondevice] co-design architectures and update rules to fit extreme SRAM budgets, often by freezing layers or pruning gradient paths. LANCE is complementary to these approaches: instead of restricting the optimization, it preserves full backpropagation while directly compressing activations. Finally, in continual learning, regularization methods [Kirkpatrick2017OvercomingNetworks, Serra2018OvercomingTask], replay-based approaches [Chaudhry2019EfficientA-GEM, Chaudhry2019OnLearning], and gradient-projection methods [Saha2021GradientLearning, Lin2022TRGP:Learning, Lin22CUBER, Saha2023ContinualProjection, apolinario2024code, SD2023] all mitigate forgetting by controlling parameter updates or storing exemplars. LANCE differs by operating directly in the activation subspace: new tasks are assigned to the orthogonal complement of previously used directions during calibration. This yields performance competitive with gradient-projection methods while requiring significantly less memory, and without rehearsal buffers.
@@ Acknowledgments
Appendix 0.A Datasets, Models and Hyperparameters
@@
For single-task fine-tuning we report top-1 accuracy on the test split, peak activation memory in MB (estimated from tensors retained for the backward pass), and training FLOPs. [^1] Peak activation memory and FLOPs formulas are detailed in Appendix 0.C. For continual learning experiments (Section 2.3), we additionally report ACC and BWT following standard CL definitions.
0.A.1 For Single-Task Fine-Tuning Experiments
0.A.1.1 Datasets
We evaluate on five image classification datasets spanning scale and difficulty: CIFAR-10 [Krizhevsky2009LearningImages], CIFAR-100 [Krizhevsky2009LearningImages], Oxford-IIIT Pets [petsdataset], Flowers-102 [flower102dataset], and CUB-200 [cub200]. We use the official training/test splits provided by each dataset and do not carve out an additional validation set. For all datasets, images are resized to match the native input resolution of the chosen architecture (e.g., $224\times 224$ for ImageNet-pretrained models), and standard data augmentation is applied consisting of random horizontal flipping.
0.A.1.2 Model Architectures
For single-task experiments we use four pretrained models: ResNet18, ResNet34 [he2016deep], MobileNetV2 [mobilenetv2], and MCUNet [Lin2020MCUNet:Devices]. All of them are initialized from ImageNet-1k pretrained weights. For each model, we unfreeze the last $N\in\{2,4\}$ convolutional layers and fine-tune them while the rest of the network remains frozen.
0.A.1.3 Training Protocol
@@ 0.A.1.4 Hyperparameters
0.A.2 For Continual Learning Experiments
@@
0.A.2.1 Model Architecture
We employ two neural network architectures in our experiments: an AlexNet-inspired model following [Serra2018OvercomingTask], and a Reduced ResNet18 as described in [Lopez-Paz2017GradientLearning]. The AlexNet-like network integrates batch normalization (BN) layers after every convolutional and fully connected layer except the final classifier. BN parameters are updated only during the first task and kept fixed thereafter. Its architecture consists of three convolutional layers with $64$, $128$, and $256$ filters, using kernel sizes of $4\times 4$, $3\times 3$, and $2\times 2$, respectively, each followed by a $2\times 2$ max-pooling operation. Two fully connected layers with $2048$ units each are appended. ReLU activations are applied throughout, and dropout is used with rates of $0.2$ for the first two layers and $0.5$ for the rest. The Reduced ResNet18 follows the setup in [Saha2021GradientLearning]. For Split miniImageNet tasks, its first convolutional layer uses stride $2$, whereas for the 5-Datasets benchmark, stride $1$ is used. Across all experiments, models are trained with cross-entropy loss.
0.A.2.2 Dataset Statistics
Tables 6 and 5 summarize the datasets employed in our continual learning benchmarks. We adopt the same train/test splits as in [Saha2021GradientLearning, Lin2022TRGP:Learning, Saha2023ContinualProjection, apolinario2024code]. For the 5-Datasets benchmark, grayscale samples are duplicated across three channels to match RGB input requirements. All images are resized to $32\times 32$, yielding an effective input dimension of $3\times 32\times 32$.
0.A.2.3 Metrics
Following standard practice in prior continual learning works [Lopez-Paz2017GradientLearning, Saha2021GradientLearning, Lin2022TRGP:Learning, Saha2023ContinualProjection, apolinario2024code], we evaluate performance using two metrics: the average final accuracy across all tasks (ACC) and the Backward Transfer (BWT), which captures the extent to which learning new tasks affects performance on previously learned ones. They are defined as:
$\text{ACC}=\sum_{i=1}^{T}\frac{{A}_{T,i}}{T}\text{ ; }\text{BWT}=\sum_{i=1}^{T-1}\frac{{A}_{T,i}-{A}_{i,i}}{T-1},$ (6)
where $T$ denotes the total number of tasks, and ${A}_{j,i}$ represents the accuracy on the $i$-th task after the model has sequentially trained up to the $j$-th task $(i\leq j)$.
0.A.2.4 Hyperparameters
@@ 0.A.3 Compute resources
Appendix 0.B Proofs for Section 2.4
@@
Since $\mathsf{P}^{(l)}$ is an orthogonal projector (Assumption A1),
@@
Thus $\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}$ is a descent direction unless it is zero.
@@
for $\eta\leq 1/L$. Summation gives $\sum_{k}\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}^{2}<\infty$ and thus $\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\to 0$. Any limit point $\bm{\theta}$$_{\star}$ satisfies $\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{\star})=\mathbf{0}$.
See 1
Proof
From Theorem˜2.2, $\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\to 0$. Then
@@
Take a limit point $\bm{\theta}$$_{\star}$.
Remark 1(Why the leakage bound is reasonable)
@@
Appendix 0.C Complexity Analysis of LANCE
@@
Appendix 0.D Additional Gradient-Compression Ablations
For completeness, we extend the gradient fidelity analysis of Sec. 3.3 to three additional fine-tuning datasets: Flowers102, Oxford-IIIT Pets, and CIFAR-10. Using the same experimental setup as in the main paper, we vary the energy threshold $\varepsilon$ to induce different activation compression ratios and measure (i) the angle between full-BP gradients and LANCE-projected gradients, (ii) the resulting memory ratio, and (iii) the accuracy obtained under each setting. Appendix Fig. 6 shows results for MCUNet and ResNet34. The behavior is highly consistent with the CUB-200 and CIFAR-100 experiments: gradient alignment degrades smoothly as compression increases, yet even aggressive compression ($\sim$2 orders of magnitude memory reduction) preserves descent directions within $70^{\circ}$. These findings demonstrate that the one-shot activation subspace generalizes well across models and datasets, further validating the stability and robustness of LANCE’s low-rank representations.
Figure 6: Additional gradient–compression ablations on Flowers, Pets, and CIFAR-10. We repeat the analysis of Fig. 4 on three additional fine-tuning datasets using MCUNet and ResNet34. For each dataset, we report the mean and standard deviation of the angle between full-BP gradients and LANCE-projected gradients over the final 10 epochs of training, together with the corresponding memory ratio (LANCE/BP) and accuracy across different energy thresholds $\varepsilon$. Across all datasets and architectures, gradient alignment degrades smoothly as $\varepsilon$ decreases, while extreme compression (up to two orders of magnitude in activation memory) continues to yield descent directions within $\sim 70^{\circ}$ of full BP. These trends closely mirror those observed in the main experiments, reinforcing the robustness and consistency of LANCE’s low-rank subspace across diverse data distributions.
Appendix 0.E Robustness of LANCE Under Small-Batch Regimes
To evaluate LANCE under more restrictive memory settings, we additionally benchmarked batch sizes $\text{BS}=4$ and $\text{BS}=8$, which are more representative of MCU- and IoT-class devices. The results (Tables 89) show that LANCE maintains the same qualitative behavior observed with $\text{BS}=128$: memory reductions of 1–2 orders of magnitude, negligible FLOPs overhead, and accuracy comparable to full BP. This indicates that LANCE’s efficiency gains do not rely on large batch sizes and extend reliably to resource-constrained edge settings.
Table 8: Performance and efficiency comparison. Results are reported in terms of accuracy, memory usage, and Giga FLOPS (GFLOPs) for different batch sizes (BS).
Mdl. Method # Layers CIFAR10 (BS = 4) CIFAR10 (BS = 8) CIFAR100 (BS = 4) CIFAR100 (BS = 8)
Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$
MCUNet BP 2 77.17 0.220 0.025 76.87 0.439 0.049 54.24 0.220 0.025 53.96 0.439 0.049
4 83.93 0.403 0.039 83.68 0.806 0.079 60.99 0.403 0.039 59.27 0.806 0.079
LANCE 2 69.72 0.007 0.014 76.32 0.011 0.028 45.70 0.008 0.014 51.27 0.010 0.028
4 74.23 0.018 0.023 77.89 0.020 0.045 49.23 0.015 0.023 53.36 0.019 0.045
ResNet34 BP 2 94.08 0.766 1.850 93.92 1.531 3.699 77.85 0.766 1.850 77.99 1.531 3.699
4 94.96 1.531 3.699 95.10 3.062 7.399 79.02 1.531 3.699 79.10 3.062 7.399
LANCE 2 91.75 0.059 1.018 92.29 0.093 2.022 73.36 0.061 1.021 74.51 0.101 2.036
4 92.05 0.168 2.114 93.20 0.232 4.124 73.09 0.156 2.096 76.17 0.252 4.161
Table 9: Performance and efficiency comparison. Results are reported in terms of accuracy, memory usage, and Giga FLOPS (GFLOPs) for different batch sizes (BS).
Mdl. Method # Layers CUB200 (BS = 4) CUB200 (BS = 8) PETS (BS = 4) PETS (BS = 8)
Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$
MCUNet BP 2 32.57 0.220 0.025 32.84 0.439 0.049 55.52 0.220 0.025 55.55 0.439 0.049
4 37.50 0.403 0.039 37.04 0.806 0.079 61.98 0.403 0.039 61.35 0.806 0.079
LANCE 2 34.79 0.011 0.015 34.97 0.013 0.029 60.32 0.011 0.015 59.85 0.015 0.029
4 35.90 0.024 0.024 36.12 0.029 0.047 60.75 0.028 0.024 59.72 0.032 0.047
ResNet34 BP 2 66.50 0.766 1.850 66.86 1.531 3.699 91.44 0.766 1.850 91.11 1.531 3.699
4 69.09 1.531 3.699 69.57 3.062 7.399 91.14 1.531 3.699 91.63 3.062 7.399
LANCE 2 65.53 0.053 1.006 64.81 0.094 2.022 91.20 0.075 1.043 90.87 0.111 2.055
4 65.72 0.160 2.102 66.31 0.220 4.102 90.84 0.182 2.136 90.92 0.275 4.202
Appendix 0.F Limitations and Future Direction.
LANCE opens several promising directions for future work. First, while the method relies on activation subspaces obtained from a small calibration set, developing mechanisms to adapt or refine these subspaces under strong domain shift or evolving representations would broaden its applicability—including extending the approach to earlier stages of pre-training, where representation drift is more pronounced. In continual learning, relaxing the strict orthogonality constraint could enable positive forward transfer between related tasks while retaining stability. Additionally, augmenting the fixed low-rank subspace with selective or lightweight updates may improve adaptability on tasks requiring larger representational changes. Finally, although LANCE is compatible with MCU-class devices, integrating it with device-specific compilation and memory-scheduling pipelines presents an exciting opportunity to realize its full hardware potential.
Appendix 0.G Code Availability
The implementation of LANCE, together with scripts to reproduce all experiments, will be released in a public repository after the review process. This will include training and evaluation code, dataset preprocessing pipelines, and experiment configuration files to ensure full reproducibility.